# Solving the Transient Dyson Equation with Quasilinear Complexity via Matrix Compression

### v1

@@ Abstract
We introduce a numerical strategy to efficiently solve the out-of-equilibrium Dyson equation in the transient regime. By discretizing the equation into a compact matrix form and applying state-of-the-art matrix compression techniques, we achieve significant improvements in computational efficiency which result in quasi-linear scaling of both time and space complexity with propagation time. This enables to compute accurate solutions even for systems with multiple and disparate time scales. We benchmark our solver by simulating a voltage-biased Josephson junction formed by a quantum dot connected to two superconducting leads.
@@ I Introduction
The Dyson equation $G=g+g\cdot\Sigma\cdot G$ is one example of such integral equation. It relates the full Green function $G$ which describes the two-points correlations in the system, to the free Green function $g$ representing the non-interacting system, and the self-energy $\Sigma$ which accounts for interaction effects. Solving this equation has proven to be challenging both at equilibrium [6] and out-of-equilibrium [7], yet some significant advancement were made by leveraging matrix compression techniques [8, 9]. In this work we refine the usage of these compressions methods to solve the out-of-equilibrium Dyson equation with unprecedented efficiency.
@@
$\small G(t,t^{\prime})=g(t,t^{\prime})+\iint_{\boldmath{-\infty}}^{\infty}g(t,t_{1})\Sigma(t_{1},t_{2})G(t_{2},t^{\prime})\text{d}t_{1}\text{d}t_{2}.$ (1)
@@
The traditional method stores each block explicitly, thus requiring $\mathcal{O}(N^{2})$ space, and setting a lower bound on the computational complexity. The equation for the retarded Dyson equation is usually solved using a time-stepping approach that requires $\mathcal{O}(N^{3})$ operations, as shown in [7, 8]. Using eq. 7 to evaluate the Keldysh component requires integrating over all the history of the system, starting at $\infty$- ∞. While there are methods to perform this integral [10, 8, 6], we impose in the following that the self-energy is turned on at $t=t_{0}$ and eventually turned off after $t_{\text{end}}$, i.e. $\forall(t,t^{\prime})\notin[t_{0},t_{\text{end}}]^{2}:\Sigma(t,t^{\prime})^{R/A/K}=0$. This simplifies the discussion by ensuring that all integration domains lie within $[t_{0},t_{\text{end}}]$.
As $N\sim\tau_{\text{long}}/\tau_{\text{short}}$, with $\tau_{\text{long}}$ and $\tau_{\text{short}}$ the longest and shortest system timescales, the traditional time-stepping method require $\mathcal{O}(\left(\tau_{\text{long}}/\tau_{\text{short}}\right)^{3})$ operations. This unfavourable scaling has prevented the use of the Dyson equation in simulating quantum system with widely separated timescales. Alternative methods that sacrifice part of the quantum field elegance and flexibility for speed have emerged [11, 11, 12]. However, recent works have reduced the required time and space to solve the Dyson equation to $\mathcal{O}(N^{2}\log(N))$ and $\mathcal{O}(N\log(N))$, respectively, by using matrix compression techniques [8]. Here, we take a step beyond by proposing a new discretization of the Dyson equation in time domain that enables a more efficient use of matrices compression method. As a result, we reduce the time complexity from a quadratic scaling to the quasilinear scaling $\mathcal{O}(N\log(N))$ without further assumptions.
@@ II Algorithm
We discretize the kernel products using the Nyström method with quadrature weights that remain constant away from the integration domain edges, a condition satisfied by Gregory quadratures [13]. For simplicity, we restrict the presentation to a low-order discretization using trapezoidal quadrature. Thus, we approximate the kernel product $A^{\text{R}}B^{\text{R}}$ as
@@
where $\odot$ is the block-wise matrix product defined as $\left[A\odot B\right]_{p,q}\equiv A_{p,q}B_{p,q}$ with $p$, $q$ block indices. $\text{Dg}\left(\mathbf{M}\right)$ is the block-diagonal matrix defined by $\text{Dg}\left(\mathbf{M}\right)_{p,q}\equiv\mathbf{M}_{p,q}\delta_{p,q}$, and $\mathbf{w}$ is the block matrix with entries: $[(1-\delta_{p,q}/2)]_{p,q}$. Similar expressions of the discretized kernel product can be derived for the other relevant products, i.e. $A^{\text{A}}B^{\text{A}}$, $A^{\text{K}}B^{\text{K}}$, $A^{\text{K}}B^{\text{R}}$ and $A^{\text{A}}B^{\text{K}}$. We can now turn to the matrix formulation of the discretized Dyson equation
@@
with $\qquad\text{with }\mathbf{S}$ $\equiv\delta_{t}\cdot\left(\frac{\text{Dg}\left(\mathbf{F^{\text{R}}}\right)}{4}-\frac{1}{2}\right)\cdot\text{Dg}\left(\mathbf{G^{\text{R}}}\right).$ (15)
@@
Table 1: Time complexity of several matrix operations. $A$ and $B$ are $N\times N$ matrices, $r$ is their maximum HSS-rank of the involved matrices. For a review of HSS algorithms, see [14].
@@
We benchmarked this algorithm by simulating the out-of-equilibrium transport in a Josephson junction formed by a quantum dot (QD) connected to two superconducting leads. The solver implementation is hosted on github [20]. All the simulation are performed on a M1-Pro CPU with 16GB of memory. The leads are superconductors described by the Barden-Cooper-Schrieffer theory [21] with a uniform real and positive order parameter $\Delta$. The operators $\psi_{\alpha,l,s}^{\dagger}$ create an electron of spin $s$ in mode $\alpha$ of lead $l$ with energy $\varepsilon_{\alpha,l}$, while $d_{s}^{\dagger}$ creates an electron of spin $s$ on the dot with energy $\varepsilon_{d}$. $t_{\alpha,l}(t)$ represents the tunnel coupling between the dot and mode $\alpha$ of lead $l$. Using the usual Nambu spinor, where $d\equiv(d_{\uparrow},d_{\downarrow}^{\dagger})^{T}$ and $\Psi_{\alpha,l}\equiv(\psi_{\alpha,l,\uparrow},\psi_{\alpha,l,\downarrow}^{\dagger})^{T}$, we can write the standard Hamiltonian as $H(t)=H_{L}+H_{D}+H_{T}(t)$ [22, 23, 24] , where
@@
[…]
@@
In the flat-band limit, and upon absorbing the density of states into the tunnelling coefficient, the retarded Green functions of the leads are
$g_{\alpha}^{R}(t)=-i\delta(t)\sigma_{0}+\Delta\left\{J_{0}(\frac{\Delta t}{\hbar})\sigma_{x}+iJ_{1}(\frac{\Delta t}{\hbar})\sigma_{0}\right\},$ (20)
@@
$g_{\text{dot}}^{\text{R}}(t)=-ie^{-i\sigma_{z}Et/\hbar}.$ (21)
@@
where $\rho_{\beta}(t,t^{\prime})\equiv-i\beta^{-1}\text{csch}\left\{\pi(t-t^{\prime})\beta^{-1}\right\}$ and $\beta\equiv 1/k_{B}T$. We name the two superconductor leads left ($L$) and right ($R$). The junction is drive out-of-equilibrium by symmetric voltage bias, i.e. $V_{L/R}(t)=\pm V/2$. The phase difference $\phi(t)=\phi_{R}(t)-\phi_{L}(t)$ satisfies $\phi(0)=0$ and we impose $\phi(t_{\text{end}})=16\pi$ by setting $V=2\hbar\phi(t_{\text{end}})/et_{\text{end}}$. We neglect any dependence of the tunnel coupling on the mode index $t_{\alpha,l}=t_{\alpha}$ and the tunnelling rate associate to the lead $l$ is $\Gamma_{l}=2|t_{l}|^{2}/\hbar$. We set $\Gamma_{\alpha}=5\Delta/\hbar$, $\beta=10^{2}\Delta^{-1}$ and $\epsilon_{d}=0$. Such Josephson junctions host a pair of discrete particle-hole symmetric bound states, known as Andreev bound states (ABS), which are entirely detached from the continuum outside the superconducting gap and exhibit avoided crossing behaviour at the Fermi energy. Each of these states carries an opposite current [25, 23].
We first measure the solver runtime for increased simulation length $t_{\text{end}}$, keeping $\delta_{t}$ constant as shown in fig. 1. At $\phi\ll 1$ the states are symmetrically occupied, and the current vanishes. However, the voltage bias induces non-adiabatic transitions that couple the ABSs to each other and to the continuum outside the superconducting gap, driving the system into a non-stationary steady state [25, 23, 26]. In the steady state, the ABSs occupation is not symmetric, hence a net current flow across the junction. The HSS-rank of the retarded Green function saturates, while the one of the kinetic Green function grows slowly enough to ensure quasi-linear solving time in $t_{\text{end}}/\delta_{t}$.
To evaluate the solver convergence rate, a high accuracy current estimate $\left\langle I_{\text{ref}}\right\rangle(t)$ is computed using a small-time step $\delta^{*}_{t}=3.9\cdot 10^{-4}\Delta/\hbar$ combined with a Richardson acceleration [27]. Writing $\left\langle I_{\delta_{t}}(t)\right\rangle$ the current evaluated for the time step $\delta_{t}$, we have:
@@
We then define the convergence error for $\delta_{t}>\delta_{t}^{*}$ as
@@ IV Conclusion
By carefully crafting a discretization of the Dyson equation and by using modern matrix compression codes, we solved the out-of-equilibrium Dyson equation in quasi-linear time and space. Benchmarks demonstrate quasilinear scaling, with the measured runtime exhibiting $\mathcal{O}(N\log N)$ behavior when using HSS compression, in contrast to the expected $\mathcal{O}(N^{3})$ trend observed in reference simulations. Utilizing a laptop, we accessed parameter regimes that are typically beyond the reach of a single-CPU machine. This quasi-linear complexity critically depends on the effectiveness of HSS compression in representing the Green functions and self-energies, which we expect to generally hold, as discussed earlier. Consequently, given the existence of scalable and efficient implementations of HSS compression or similar methods, and considering their robustness, we anticipate that this solving technique will be capable of addressing extremely large problems, provided that the compressed representations of the kernels fit in memory.

### v2

@@ Abstract
We introduce a numerical strategy to efficiently solve the out-of-equilibrium Dyson equation in the transient regime. By discretizing the equation into a compact matrix form and applying state-of-the-art matrix compression techniques, we achieve significant improvements in computational efficiency which results in quasi-linear scaling of both time and space complexity with propagation time. This enables to compute accurate solutions even for systems with multiple and disparate time scales. We benchmark our solver by simulating a voltage-biased Josephson junction formed by a quantum dot connected to two superconducting leads.
@@ I Introduction
The Dyson equation $G=g+g\cdot\Sigma\cdot G$ is one example of such integral equation. It relates the full Green function $G$ which describes the two-points correlations in the system, to the free Green function $g$ representing the non-interacting system, and the self-energy $\Sigma$ which accounts for interaction effects. Solving this equation has proven to be challenging both at equilibrium [6] and out-of-equilibrium [7], yet some significant advancements were made by leveraging matrix compression techniques [8, 9]. In this work, we refine the usage of these compressions methods to solve the out-of-equilibrium Dyson equation with unprecedented efficiency.
@@
$\small G(t,t^{\prime})=g(t,t^{\prime})+\iint_{{-\infty}}^{\infty}g(t,t_{1})\Sigma(t_{1},t_{2})G(t_{2},t^{\prime})\text{d}t_{1}\text{d}t_{2}.$ (1)
@@
The traditional method stores each block explicitly, thus requiring $\mathcal{O}(N^{2})$ space, and setting a lower bound on the computational complexity. The retarded Dyson equation is usually solved using a time-stepping approach demanding $\mathcal{O}(N^{3})$ operations as $N$ steps must be computed for $N$ values of $t^{\prime}$ and each step requires summing over the full system history to evaluate the integral as shown in [7, 8]. Using eq. 7 to evaluate the Keldysh component requires integrating over all the history of the system, starting at -$\infty$. While there are methods to perform this integral [10, 8, 6], we impose in the following that the self-energy is turned on at $t=t_{0}$ and eventually turned off after $t_{\text{end}}$, i.e. $\forall(t,t^{\prime})\notin[t_{0},t_{\text{end}}]^{2}:\Sigma(t,t^{\prime})^{R/A/K}=0$. This simplifies the discussion by ensuring that all integration domains lie within $t_{0},t_{\text{end}}]$.
As $N\sim\tau_{\text{long}}/\tau_{\text{short}}$, with $\tau_{\text{long}}$ and $\tau_{\text{short}}$ the longest and shortest system timescales, the traditional time-stepping method require $\mathcal{O}(\left(\tau_{\text{long}}/\tau_{\text{short}}\right)^{3})$ operations. This unfavourable scaling has prevented the use of the Dyson equation in simulating quantum system with widely separated timescales. Alternative methods that sacrifice part of the quantum field elegance and flexibility for speed have emerged [11, 11, 12], or solver specialized for specific form of the equation [13, 14, 9]. However, recent works have reduced the required time and space to solve the Dyson equation to $\mathcal{O}(N^{2}\log(N))$ and $\mathcal{O}(N\log(N))$, respectively, by using matrix compression techniques to perform fast integral evaluation and compress the kernel representation [8]. Here, we take a step beyond by proposing a new discretization of the Dyson equation in time domain that enables a more efficient use of matrices compression method. As a result, we reduce the time complexity from a quadratic scaling to the quasilinear scaling $\mathcal{O}(N\log(N))$ without further assumptions.
@@ II Algorithm
We discretize the kernel products using the Nyström method with quadrature weights that remain constant away from the integration domain edges, a condition satisfied by Gregory quadratures [15]. To avoid overwhelming the reader with indices, we propose here a compact formulation of the algorithm and we restrict the presentation to a low-order discretization using trapezoidal quadrature. An expanded version is presented in the supplemental materials [16]. Thus, we approximate the kernel product $A^{\text{R}}B^{\text{R}}$ as
@@
where $\odot$ is the block-wise matrix product defined as $\left[A\odot B\right]_{p,q}\equiv A_{p,q}B_{p,q}$ with $p$, $q$ block indices. $\text{Dg}\left(\mathbf{M}\right)$ is the block-diagonal matrix defined by $\text{Dg}\left(\mathbf{M}\right)_{p,q}\equiv\mathbf{M}_{p,q}\delta_{p,q}$, and $\mathbf{w}$ is the block matrix with entries: $(1-\delta_{p,q}/2)]_{p,q}$. Similar expressions of the discretized kernel product can be derived for the other relevant products, i.e. $A^{\text{A}}B^{\text{A}}$, $A^{\text{K}}B^{\text{K}}$, $A^{\text{K}}B^{\text{R}}$ and $A^{\text{A}}B^{\text{K}}$. We can now turn to the matrix formulation of the discretized Dyson equation
@@
$\qquad\text{with }\mathbf{S}$ $\equiv\delta_{t}\cdot\left(\frac{\text{Dg}\left(\mathbf{F^{\text{R}}}\right)}{4}-\frac{1}{2}\right)\cdot\text{Dg}\left(\mathbf{G^{\text{R}}}\right).$ (15)
@@
Table 1: Time complexity of several matrix operations. AA and $B$ are $N\times N$ matrices, $r$ is their maximum HSS-rank of the involved matrices. For a review of HSS algorithms, see [17].
@@
We benchmarked this algorithm by simulating the out-of-equilibrium transport in a Josephson junction formed by a quantum dot (QD) connected to two superconducting leads as this system is well known [23, 24, 25, 26]. The solver implementation is hosted on github [27]. All the simulations were performed on a ultra 265K CPU with 48GB of RAM. The leads are superconductors described by the Barden-Cooper-Schrieffer theory [28] with a uniform real and positive order parameter $\Delta$. The operators $\psi_{\alpha,l,s}^{\dagger}$ create an electron of spin $s$ in mode $\alpha$ of lead $l$ with energy $\varepsilon_{\alpha,l}$, while $d_{s}^{\dagger}$ creates an electron of spin $s$ on the dot with energy $\varepsilon_{d}$. $t_{\alpha,l}(t)$ represents the tunnel coupling between the dot and mode $\alpha$ of lead $l$. Using the usual Nambu spinor, where $d\equiv(d_{\uparrow},d_{\downarrow}^{\dagger})^{T}$ and $\Psi_{\alpha,l}\equiv(\psi_{\alpha,l,\uparrow},\psi_{\alpha,l,\downarrow}^{\dagger})^{T}$, we can write the standard Hamiltonian as $H(t)=H_{L}+H_{D}+H_{T}(t)$ [24, 25, 26] , where
@@
$H_{T}(t)$ $=\sum_{\alpha,l}\left\{\Psi_{\alpha,l}^{\dagger}\mathcal{T}_{\alpha,l}(t)d+\text{H.c}\right\},$ (18)
with $\mathcal{T}_{\alpha,l}\equiv t_{\alpha,l}\sigma_{z}e^{i\sigma_{z}\phi_{\alpha,l}(t)/2}$ being the time-dependent tunnel matrix, $\sigma_{i}$ the Pauli matrices acting in the Nambu space, the transmission phase $\phi_{l}(t)\equiv\frac{2e}{\hbar}\int^{t}{V_{l}(\tau)}d\tau$, e-e the charge of an electron and $V_{l}(\tau)$ the voltage bias from the QD to the lead $l$. Employing the quantum field formalism [2, 1], we integrate out the infinite leads, resulting in an effective description of the dot in terms of its bare Green function $g_{d}$ dressed by the self-energies $\Sigma_{l}$ induced by the leads
@@
In the flat-band limit, and upon absorbing the density of states into the tunneling coefficient, the retarded Green functions of the leads are
$g_{\alpha}^{R}(t)=-i\pi\delta(t)\sigma_{0}+\pi\Delta\left\{J_{0}(\frac{\Delta t}{\hbar})\sigma_{x}+iJ_{1}(\frac{\Delta t}{\hbar})\sigma_{0}\right\},$ (20)
@@
$g_{\text{dot}}^{\text{R}}(t)=-ie^{-i\sigma_{z}\varepsilon_{d}t/\hbar}.$ (21)
@@
where $\rho_{\beta}(t,t^{\prime})\equiv-i\beta^{-1}\text{csch}\left\{\pi(t-t^{\prime})\beta^{-1}\right\}$ and $\beta\equiv 1/k_{B}T$. We name the two superconductor leads left ($L$) and right ($R$). For $t\geq 0$, the junction is drive out-of-equilibrium by symmetric voltage bias, i.e. $V_{L/R}(t)=\pm V/2$. The phase difference $\phi(t)=\phi_{R}(t)-\phi_{L}(t)$ satisfies $\phi(0)=0$. We neglect any dependence of the tunnel coupling on the mode index $t_{\alpha,l}=t_{\alpha}$ and the tunnelling rate associate to the lead $l$ is $\Gamma_{l}=\pi|t_{l}|^{2}/\hbar$. Such Josephson junctions host a pair of discrete particle-hole symmetric bound states, known as Andreev bound states (ABS), which are entirely detached from the continuum outside the superconducting gap and exhibit avoided crossing behaviour at the Fermi energy when the dot is not resonant. Each of these states carries an opposite current [29, 25]. We set $\Gamma_{\alpha}=5\Delta/\hbar$, $\beta=10^{2}\Delta^{-1}$ and $\epsilon_{d}=0$.
For the first set of experiments, we impose $\phi(t_{\text{end}})=16\pi$ by setting $V=2\hbar\phi(t_{\text{end}})/et_{\text{end}}$. Thus, we have $\phi(t)=16\pi t/t_{\text{end}}$. We measure the solver runtime for increased simulation length $t_{\text{end}}$, keeping $\delta_{t}$ constant as shown in fig. 1. At $\phi\ll 1$ the states are symmetrically occupied, and the current vanishes. However, the voltage bias induces non-adiabatic transitions that couple the ABSs to each other and to the continuum outside the superconducting gap, driving the system into a non-stationary steady state [29, 25, 30]. In the steady state, the ABSs occupation is not symmetric, hence a net current flow across the junction. The HSS-rank of the retarded Green function saturates, while the one of the kinetic Green function grows slowly enough to ensure quasi-linear solving time in $t_{\text{end}}/\delta_{t}$.
The simulation error arises from the interplay between discretization error and compression error. When the compression tolerance is sufficiently small, the error is dominated by the discretization error $\epsilon(\delta_{t})$, which can be assumed to have the form
$\epsilon(\delta_{t})=\sum_{p\geq\alpha+1}a_{p}\,\delta_{t}^{p}+o(\delta_{t}^{n}),$ (25)
where $\alpha\in\mathbb{N}^{+}$ is the convergence rate. By performing $n_{\text{sim}}$ simulations with varying time steps $\delta_{t,0<i\leq n_{\text{sim}}}$, one can evaluate $n_{\text{sim}}-1$ terms of this expansion and construct a simulation result whose discretization error $\epsilon(\delta_{t,1},...,\delta_{t,n})$ is of the form
$\epsilon(\delta_{t,1},...,\delta_{t,n})=\sum_{p\geq\alpha+n_{\text{sim}}}a_{p}^{\prime}\delta_{t}^{p}+o(\delta_{t}^{n}),$ (26)
Hence, this increases the method convergence rate from $\alpha$ to at least $\alpha+n_{\text{sim}}$. This classical method is known as Richardson acceleration [31]. To evaluate the solver convergence rate, we first compute a high-accuracy reference current estimate $\left\langle I_{\text{ref}}\right\rangle(t)$ using Richardson acceleration
@@
where $\delta^{*}_{t}=3.9\cdot 10^{-4}\Delta/\hbar$ and $\left\langle I_{\delta_{t}}(t)\right\rangle$ the current evaluated for the time step $\delta_{t}$. By construction, this estimate as a convergence rate of at least $3$ in $\delta^{*}_{t}$. We then define the convergence error for $\delta_{t}>\delta_{t}^{*}$ as
@@
Figure 3: I-V characteristics of a voltage-biased quantum-dot Josephson junction with symmetric coupling $\Gamma=\Gamma_{L}=\Gamma_{R}$ at resonance $\varepsilon=0$. Transient simulations run from $t=0$ to $t_{\text{end}}=\max(200h/\Delta,20T_{\text{J}})$ with compression tolerances $\epsilon_{\text{rel}}=\epsilon_{\text{abs}}=10^{-6}$ in simulation units. Results reproduce the classical behavior from. [25].
We extract the low-frequency averaged current, $I_{\text{LF}}(t)$, by averaging the instantaneous current $\langle I_{l}(t)\rangle$ over two Josephson periods, $2T_{\text{J}}=2\pi/(eV\hbar)$, accounting for the fractional Josephson radiation characteristic of a quantum dot Josephson junction at resonance [30]
$I^{\text{LF}}_{\delta_{t}}(t)\equiv\frac{1}{2T_{\text{J}}}\int_{t-2T_{\text{J}}}^{t}\langle I_{\delta_{t}}(t^{\prime})\rangle\,dt^{\prime}.$ (29)
The system is simulated from to $t_{0}=0$ to $t_{\text{end}}=\max(200\,\hbar/\Delta,\ 20\,T_{\text{J}})$ to accommodate the different timescales across voltage regimes. At low voltage, the longest timescale is the Josephson period $T_{\text{J}}$, where interactions between the ABSs and the continuum occur primarily at anti-crossings, requiring simulation over many Josephson periods. At higher bias, however, the stronger couplings allow the system to reach steady state more rapidly, eliminating the need for such extended simulations. We use an adaptive multi-step Richardson extrapolation to both increase the result accuracy and provide an error estimate. The implementation is part of the solver module see [27]. The step size is refined, until the absolute or relative error estimate are below $10^{-2}$ in simulation units. The initial time step is set to $\min(0.3T_{\text{J}},0.5/\Gamma)$ to ensure proper resolution of the fast timescale.
As the ABSs are separated from the continuum by a finite gap $\Delta_{\text{ABS}}$. For $V\ll f(\Delta_{\text{ABS}})$, the ABSs become effectively decoupled from the continuum [30, 29]. In the absence of relaxation mechanisms, their long-time occupations depend on initial conditions, implying no well-defined steady state exists. Nevertheless, $I_{\text{LF}}(t)$ rapidly approaches a stationary regime, enabling extraction of meaningful current-voltage characteristics even at low voltages. This behavior is illustrated in fig. 3, which reproduces the classical results from [25].
@@ IV Conclusion
By carefully crafting a discretization of the Dyson equation and by using modern matrix compression codes, we solved the out-of-equilibrium Dyson equation in quasi-linear time and space. Benchmarks demonstrate quasilinear scaling, with the measured runtime exhibiting $\mathcal{O}(N\log N)$ behavior when using HSS compression, in contrast to the expected $\mathcal{O}(N^{3})$ trend observed in reference simulations. Utilizing a desktop, we accessed parameter regimes that are typically beyond the reach of a single-CPU machine. This quasi-linear complexity critically depends on the effectiveness of HSS compression in representing the Green functions and self-energies, which we expect to generally hold, as discussed earlier. Consequently, given the existence of scalable and efficient implementations of HSS compression or similar methods, and considering their robustness, we anticipate that this solving technique will be capable of addressing extremely large problems, provided that the compressed representations of the kernels fit in memory.